Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Difficult Math Problem #81 - Probability

Expert replies
by 800guy » Fri Dec 29, 2006 10:06 am
From the even numbers between 1 and 9, two different even numbers are to be chosen at random. What is the probability that their sum will be 8?

answer coming when a few people answer.
Join the discussion
Source: — Problem Solving |

by thankont » Fri Dec 29, 2006 2:51 pm
My ans. is 1/4C2 = 1/6 (one pair will do the job (2,6) if of course we assume that we cannot pick same number twice.
Join the discussion

by maxim730 » Fri Dec 29, 2006 7:21 pm
1/3?
Join the discussion

by vidskris » Sat Dec 30, 2006 10:20 am
THE ONLY WAY WHEN SUM IS 8 IS WEN THE 2 NUMBERS ARE 2, 6
SO IT SHOULD BE 1/10 x 2/9= 1/45
Join the discussion

by miav » Sun Dec 31, 2006 6:22 am
There are 4 even numbers. So the all the possible set is 1/4 * 1/4.
The only way to 2 get sum of 8 is (2,6) and (6,2). Here these are 2 different set and should be counted twice.
So that answer should be 2/16 = 1/8.
Join the discussion

Answer

by gloubiboulga » Sun Dec 31, 2006 10:17 am
There are 4 even integers: 2,4,6,8.
The probability to pick 2 or 6 = 1/4 + 1/4 = 1/2
Then the probability to pick the other integer (either 2 or 6) is 1/3
So the answer is 1/2 * 1/3 = 1/6
Join the discussion

by maxim730 » Sun Dec 31, 2006 11:01 am
what's the OA? THere are like 5 different answers here :lol:
Join the discussion

OA

by 800guy » Sun Dec 31, 2006 12:28 pm
here's the OA:

Initially you have 4 even numbers (2,4,6,8 )
you can get the sum of 8 in two ways => 2 + 6 or 6 + 2
so the first time you pick a number you can pick either 2 or 8 - a total of 2 choices out of 8 => 1/2
after you have picked your first number and since you have already picked 1 number you are left with only 2 options => either (4,6,8 ) or (2,4,8 ) and you have to pick either 6 from the first set or 2 from the second and the probability of this is 1/3. Since these two events have to happen together we multiply them. ½ * 1/3 = 1/6
Join the discussion

by aim-wsc » Tue Jan 02, 2007 8:21 am
that google smiley 8) looks great :lol:
nice explanation :) !

dear 800guy
would you provide explanation for question no. 71 too?
Join the discussion

by sfyn2it » Tue Jan 17, 2012 7:09 pm
You have four choices for even numbers (2, 4, 6 and 8)

Two of these numbers need to add to 8. Only one combination allows that (2 and 6).

There are two different ways to select 2 and 6. Either select 2 in your first attempt and select 6 in your second or vise versa. The probability will be the same.

Lets take the first scenario and figure out the probability.

First for us to get a 2 from 4 numbers (2,4,6 and 8), the probability is 1/4

Now we have 3 numbers left (4, 6 and 8). To get a 6 from these 3 numbers, the probability is 1/3

Multiply the two probabilities:

1/4*1/3 = 1/12

Now since the probability of the second scenario (where we pick 6 first and 2 second) is the same, you can add the two probabilities together:

1/12 + 1/12 = 2/12 = 1/6
Join the discussion

by ronnie1985 » Tue Jan 17, 2012 9:23 pm
HERE THE EVENT SUM =8 CAN BE OBTAINED BY THE COMBINATION OF THE FOLLOWING EVENTS:-
EVENT=2 AND EVENT =6 OR EVENT =6 AND EVENT =2
P(SUM=8) = 1/4*1/3+1/4*1/3 = 1/12+1/12 = 1/6
Follow your passion, Success as perceived by others shall follow you
Join the discussion

by LalaB » Tue Jan 17, 2012 9:33 pm
2 4 6 8 are even numbers between 1 and 9

2+6 is the required sum


so, 1/4C2=1/6
Join the discussion

by jayoptimist » Wed Jan 18, 2012 1:38 am
(2,4,6,8) are the even numbers betwen 1 and 9.
we can choose a sum of 8 in 2 ways, a) 2+6 OR b)6+2.
a)
1) Probability of getting a '2' out of 4 numbers = 1/4
2) Probabiltiy of getting a '6' out of 3 numbers = 1/3
Since we want both (1) AND (2) , 1/4*1/3 = 1/12

b)
The same values result as above , i.e., 1/12

Since our requirement is either a) or b) 1/12 + 1/12 = 2/12 = 1/6.
Join the discussion

by sam2304 » Thu May 24, 2012 11:18 pm
We can get the sum in two ways 2+6 or 6+2.
Total outcomes = 4C2

So probability of getting a 8 as the sum = 2/4C2 = 2/6 = 1/3

Where am i going wrong ? ? Can someone explain me ?

Should we use permutation here while calculating the total outcomes as 2+6 and 6+2 are considered two different options ? ?
Getting defeated is just a temporary notion, giving it up is what makes it permanent.
https://gmatandbeyond.blogspot.in/
Join the discussion

by aneesh.kg » Sat May 26, 2012 2:46 am
sam2304 wrote:We can get the sum in two ways 2+6 or 6+2.
Total outcomes = 4C2

So probability of getting a 8 as the sum = 2/4C2 = 2/6 = 1/3

Where am i going wrong ? ? Can someone explain me ?

Should we use permutation here while calculating the total outcomes as 2+6 and 6+2 are considered two different options ? ?
Hi sam2304,

Let me review a few basics of Permutations & Combinations for you before I explain your doubt.

(i) Number of ways of arranging/distributing 'n' different objects on 'n' places is given by n!

(ii) Number of ways of selecting/choosing/picking 'r' objects from 'n' different objects is given by nCr. The 'C' stands for 'Combinations'.
For example:
If there are 4 people: A, B, C and D, the following are the various ways of choosing a 2-member team out of them.
(A, B), (B, C), (C, D), (A, C), (B, D), (A, D). There are 4C2 = 6 such teams. Notice that we are just concerned with the members in the team and not with arranging them.

(iii) Number of ways of arranging/distributing 'n' different objects on 'r' places (where r < n) is given by nPr. The 'P' stands for 'Permutations'.
For example:
If there are 4 people: A, B, C and D, the following are the various ways of distributing Ist and IInd prizes to 2 of them:
(A 1st, B 2nd), (B 1st, A 2nd), (B 1st, C 2nd), (C 1st, B 2nd), (C 1st, D 2nd), (D 1st, C 2nd) (A 1st, C 2nd), (C 1st, A 2nd), (B 1st, D 2nd), (D 1st, B 2nd), (A 1st, D 2nd), (D 1st, A 2nd). There are 4P2 = 12 such ways. Notice that we were concerned with arranging the two people.

(iv) OR is Addition, AND is addition.

Alright then, Do you see your mistake now?

The problem uses the word 'chosen', which indicates that we are not concerned with the arrangement. The 4C2, which is the total number of outcomes, is the number of ways of selecting 2 even numbers from 4 even numbers. It includes these: (2,4), (4,6), (6,8), (2,6), (4,8), (2,8).

How many of these add up to 8?

Just one: (2,6).

Required Probability = 1/6

An interesting Permutations & Combinations concept:
https://www.beatthegmat.com/important-pe ... 11283.html
Aneesh Bangia
GMAT Math Coach
[email protected]

GMATPad:
Facebook Page: https://www.facebook.com/GMATPad
Join the discussion